Circular No. 7286

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999ej IN NGC 495 A. Friedman, J. Y. King, and W. D. Li, University of California at

Berkeley (UCB), report the discovery by the Lick Observatory Supernova Search (LOSS; cf. IAUC 6627, 7126) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova in NGC 495. SN 1999ej was discovered and confirmed on unfiltered images taken on Oct. 18.3 and 19.3 UT (both with mag ∼ 18.1). The new object is located at α = 1^{h}22^{m}57^{s}.38, δ = +33^{o}27′58′′.0 (equinox 2000.0), which is 17′′.7 east and 20′′.1 south of the nucleus of NGC 495. A KAIT image of the same field on Oct. 15.3 (limiting mag ∼ 19.0) showed nothing at this position.

SUPERNOVA 1999ek in UGC 3329 R. Johnson and W. D. Li, UCB, also report the discovery by the LOSS

with the KAIT of an apparent supernova in UGC 3329. SN 1999ek was discovered and confirmed on unfiltered images taken on Oct. 20.5 (mag ∼ 18.1) and 21.4 UT (mag ∼ 17.9). The new object is located at α = 5^{h}36^{m}31^{s}.60, δ = +16^{o}38′17′′.8 (equinox 2000.0), which is 11′′.7 west and 12′′.1 south of the nucleus of UGC 3329. A KAIT image of the same field on 1998 Oct. 21.5 (limiting mag ∼ 19.5) showed nothing at this position.

NOVA IN SMALL MAGELLANIC CLOUD 1999 No. 3 J.-F. Glicenstein, Département d’Astrophysique, de Physique des Par-

ticules, de Physique Nucléaire et d’Instrumentation Associée, Saclay, on behalf of the EROS2 collaboration, reports the discovery of an apparent nova that triggered the EROS2 microlensing alert system. The new object is located at α = 0^{h}59^{m}23^{s}.0, δ = −73^{o}07′56′′ (equinox 2000.0), and it peaked in brightness some time between consecutive observations on Sept. 17.1 and 18.2 UT. Prior to eruption, on Sept. 17.1, the closest star (within 2′′) had an apparent magnitude of R_{c} = 19.6; the limiting magnitude on the exposures is roughly 20 in both the EROS ‘blue’ and ‘R’ bands. On Sept. 18.2, the nova was at R_{c} = 13.3; it reached a maximum around Sept. 22, saturating the CCDs (R_{c} < 12.8; the EROS R-band flux increased by a factor of 3 compared to that on Sept. 18). It is now fading; recent R_{c} magnitudes: Oct. 17, 13.3; 18, 13.7, 19, 13.9. See finding charts at http://www-dapnia.

cea.fr/Phys/Spp/Experiences/EROS/novae/en-smc-99-2/index.html.

1999 October 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green